### v2

# Attention Isn’t All You Need for Emotion Recognition: Domain Features Outperform Transformers on the EAV Dataset

Abstract
We present a systematic study of multimodal emotion recognition using the EAV dataset, investigating whether complex attention mechanisms improve performance on small datasets. We implement three model categories: baseline transformers (M1), novel factorized attention mechanisms (M2), and improved CNN baselines (M3). Our experiments show that sophisticated attention mechanisms consistently underperform on small datasets. M2 models achieved 5 to 13 percentage points below baselines due to overfitting and destruction of pretrained features. In contrast, simple domain-appropriate modifications proved effective: adding delta MFCCs to the audio CNN improved accuracy from 61.9% to 65.56% (+3.66pp), while frequency-domain features for EEG achieved 67.62% (+7.62pp over the paper baseline). Our vision transformer baseline (M1) reached 75.30%, exceeding the paper’s ViViT result (74.5%) through domain-specific pretraining, and vision delta features achieved 72.68% (+1.28pp over the paper CNN). These findings demonstrate that for small-scale emotion recognition, domain knowledge and proper implementation outperform architectural complexity.
Attention Isn’t All You Need for Emotion Recognition:
Domain Features Outperform Transformers on the EAV Dataset
Anmol Guragain Universidad Politécnica de Madrid E.T.S. Ingenieros de Telecomunicación anmol.g@upm.es
1 Introduction
Emotion recognition from physiological and behavioral signals is a core challenge in affective computing, with applications in human-computer interaction, mental health monitoring, and assistive technologies. The EAV dataset Kim et al. (2024) provides synchronized EEG, Audio, and Video recordings from 42 participants engaged in conversational interactions, offering a unique resource for multimodal emotion research alongside large-scale multimodal affect datasets such as CMU-MOSEI Bagher Zadeh et al. (2018). However, with only approximately 280 training samples per subject after data splitting, this dataset presents significant challenges for deep learning approaches.
This limited data availability motivates our central research question: Do complex attention mechanisms improve emotion recognition performance on small datasets, or do simpler approaches with domain-appropriate features perform better?
To answer this question, we systematically evaluate three categories of models. First, we implement baseline transformer architectures (M1) to establish reference performance. Second, we develop novel factorized attention mechanisms (M2) tailored to each modality’s unique structure: spatial-temporal-asymmetry attention for EEG, temporal-frequency dual attention for audio, and ViViT-style space-time attention for video. Third, we explore minimal CNN improvements (M3) focusing on bug fixes and domain-appropriate feature engineering.
Our study makes three contributions: (1) we show that factorized attention mechanisms consistently fail on small datasets, underperforming baselines by 5 to 13 percentage points; (2) we demonstrate that simple, domain-appropriate modifications (delta MFCCs for audio and band power features for EEG) achieve state-of-the-art results on this dataset; and (3) we provide analysis of failure modes and success factors for emotion recognition on limited data.
2 Related Work
2.1 Emotion Recognition Datasets
Traditional EEG datasets such as DEAP Koelstra et al. (2012) and MAHNOB-HCI Soleymani et al. (2012) employ passive elicitation paradigms where participants watch emotional stimuli. Audio-visual datasets like IEMOCAP Busso et al. (2008) capture conversational emotions but lack neurophysiological signals. The EAV dataset bridges this gap by providing synchronized multimodal recordings during active conversation, though its relatively small scale (42 participants) constrains the complexity of applicable models.
2.2 Transformer Architectures for Emotion Recognition
The Audio Spectrogram Transformer (AST) Gong et al. (2021) treats spectrograms as images, applying Vision Transformer architecture with strong performance on large-scale audio datasets. Video Vision Transformer (ViViT) Arnab et al. (2021) extends this approach to video through factorized space-time attention, achieving 22 computational efficiency over full attention. EEGformer Hussain and Park (2022) adapts transformers for EEG signals. These architectures achieve impressive results on large datasets, but their effectiveness on small-scale emotion recognition remains unclear.
2.3 Frontal Asymmetry in Emotion
Davidson’s frontal asymmetry model Davidson (1992) establishes that left frontal cortex activation (measured at F3) correlates with approach motivation and positive emotions, while right frontal activation (F4) correlates with withdrawal motivation and negative emotions. This neurophysiological insight motivates our M2 EEG architecture and M3 band power features.
2.4 Feature Engineering for Emotion Recognition
Classical approaches to emotion recognition rely on hand-crafted features. For EEG, band power and differential entropy features computed from frequency bands (delta, theta, alpha, beta, gamma) have proven effective on the SEED dataset Zheng and Lu (2015) and remain widely used across EEG emotion recognition studies Patel et al. (2021). For audio, Mel-frequency cepstral coefficients (MFCCs) and their temporal derivatives (delta and delta-delta) capture both spectral content and dynamics Schuller and others (2013); El Ayadi et al. (2011); Ververidis and Kotropoulos (2006), and are commonly extracted using toolkits such as openSMILE Eyben et al. (2010). Squeeze-and-Excitation networks Hu et al. (2018) introduced channel attention for CNNs, learning to weight feature map importance through a bottleneck architecture.
3 Dataset
The EAV dataset contains recordings from 42 participants, each providing 200 interaction samples across five emotion classes: Neutral, Anger, Happiness, Sadness, and Calmness. EEG signals are recorded using a 30-channel BrainAmp system at 500 Hz, audio at 16 kHz, and video at 30 fps.
Following standard preprocessing, EEG signals are downsampled to 100 Hz and bandpass filtered (0.5 to 45 Hz), yielding segments of shape (30 channels, 500 samples). Audio clips of 5 seconds are extracted, and 25 video frames are sampled per segment. The data is split 70/30 for training and testing per subject, resulting in approximately 280 training and 120 test samples per subject per modality.
This limited training data (280 samples with 15,000 EEG features, 80,000 audio samples, or 252242243 video pixels) fundamentally constrains model complexity and motivates our investigation of simpler approaches.
4 Methods
4.1 M1: Baseline Transformer Architectures
Our baseline models replicate the paper’s transformer approach using pretrained architectures from Hugging Face.
EEG Transformer.
We implement a custom architecture with convolutional feature extraction followed by transformer encoding. Two 1D convolutional layers (kernel size 11, padding 5) transform the 30-channel input into 60-dimensional features while preserving temporal resolution. Six transformer encoder layers with 4 attention heads and hidden dimension 60 model temporal dependencies. Mean pooling followed by a two-layer classifier produces emotion predictions. This architecture applies standard self-attention across time, treating channels as input features rather than a separate dimension.
Audio Spectrogram Transformer.
We fine-tune the pretrained AST model (MIT/ast-finetuned-audioset) with a custom 5-class classifier head. The model converts audio waveforms to mel spectrograms (1024128), extracts 1616 patches yielding approximately 1212 tokens, and applies 12 transformer layers with 768-dimensional embeddings. Training proceeds in two stages: 10 epochs with frozen backbone (learning rate 5e-4), followed by 15 epochs with unfrozen backbone (learning rate 5e-6).
Vision Transformer.
We employ a ViT model pretrained on facial emotion recognition (dima806/facial_emotions_image_detection). Each video is processed as 25 independent frames, with per-frame predictions averaged for the final classification. The same two-stage training strategy is applied: 10 epochs frozen, then 5 epochs unfrozen. While this approach ignores temporal dynamics across frames, the facial emotion pretraining (often built from large-scale facial affect data such as AffectNet) provides strong baseline performance Mollahosseini et al. (2017).
4.2 M2: Factorized Attention Mechanisms
Our main architectural contribution explores whether factorized attention, tailored to each modality’s structure, can improve upon baselines. Figure 1 illustrates the three M2 architectures.
4.2.1 M2-EEG: Tri-Stream Transformer
EEG signals contain information along three distinct axes. The spatial dimension captures relationships between brain regions (channels), the temporal dimension captures signal dynamics, and the hemispheric asymmetry dimension captures left-right differences relevant for emotional valence.
Our architecture first applies convolutional feature extraction, then branches into three parallel attention streams. The spatial stream transposes features to (time, channels, features) and applies attention across channels, learning which brain regions interact at each time point. The temporal stream applies attention across time points for each channel, capturing event-related dynamics. The asymmetry stream computes difference signals between six hemisphere pairs (Fp1-Fp2, F3-F4, F7-F8, C3-C4, P3-P4, O1-O2) based on the 10-20 electrode system, then applies attention across these pairs. Following Davidson’s model, F3-F4 receives a 1.2 attention weight.
Stream outputs are fused using learned softmax weights, with a learnable skip connection preserving convolutional features. The skip weight is initialized such that approximately 73% of the signal comes from the convolutional path. We explored three model sizes (138K, 419K, and 1.6M parameters), with the smallest recommended for this dataset size.
4.2.2 M2-EEG: EEGNet-Improved
As a literature-based alternative, we enhanced EEGNet with techniques from recent papers: squeeze-excitation blocks for channel attention, multi-scale temporal convolution (kernels 16, 32, 64), residual connections, data augmentation (amplitude scaling, Gaussian noise, time shift), label smoothing (0.1), AdamW optimizer with weight decay, and cosine annealing learning rate schedule.
4.2.3 M2-Audio: Temporal-Frequency Dual Attention
Audio spectrograms have natural 2D structure: the temporal axis captures speech rhythm, pauses, and intensity changes, while the frequency axis captures pitch, harmonics, and voice quality. Standard AST flattens this structure into a 1D patch sequence, losing the distinction.
Our architecture preserves AST’s pretrained features but adds factorized attention. After extracting 1212 patches from AST, we reshape them into a 10112 grid (timefrequency). Learnable 2D positional encodings are added. Two dual-attention blocks apply temporal attention (each frequency band attends across time) followed by frequency attention (each time step attends across frequencies). A learnable skip connection, initialized with weight (yielding ), blends the attention output with AST’s original CLS token, preserving approximately 88% of pretrained features.
4.2.4 M2-Vision: Factorized Space-Time Attention
Video understanding requires both spatial analysis within frames and temporal modeling across frames. Full space-time attention over 25 frames 196 patches has prohibitive million attention entries.
Following ViViT’s Model 2 (factorized encoder), we decompose this into spatial attention within each frame () followed by temporal attention for each patch position across frames (), achieving 22 speedup. Four factorized attention blocks process ViT’s patch embeddings (shape: batch25196768), with separate learnable spatial and temporal positional encodings. A skip connection preserves ViT’s pretrained features.
4.3 M3: CNN Improvements
After M2’s poor results, we investigated whether simple fixes to baseline CNNs could outperform complex attention mechanisms.
4.3.1 M3-Audio: Delta MFCCs
The original audio CNN uses 180 features (40 MFCCs, 12 chroma, 128 mel), averaged over time. To introduce temporal dynamics with minimal additional capacity, we augment the feature set with delta MFCCs. Delta MFCCs approximate the first temporal derivative of the cepstral coefficients and capture short-term spectral change patterns that are informative for affective speech (e.g., variations in pitch and energy).
Our modification adds 40 delta MFCC features (also time-averaged), increasing the feature vector from 180 to 220 dimensions. Combined with AdamW optimizer (weight decay 1e-4) and label smoothing (0.1), this minimal change achieved our best audio result.
4.3.2 M3-EEG v1: Bug Fixes
We identified an implementation issue in the original EEGNet classifier: applying nn.Softmax(dim=1) before CrossEntropyLoss, which already applies (log-)softmax internally. This double normalization can distort gradients and hinder optimization. We therefore removed the explicit softmax and trained the model on raw logits, using AdamW with weight decay (0.01) and label smoothing (0.1).
4.3.3 M3-EEG v2: Band Power Features
Bug fixes alone proved insufficient, as raw time-domain EEG is often heavily contaminated by artifacts (e.g., muscle activity, eye blinks, and environmental noise). Following established practice in EEG emotion recognition (e.g., SEED-style pipelines) Zheng and Lu (2015), we therefore extracted frequency-domain features.
Band power features compute power spectral density using Welch’s method for five frequency bands (delta: 0.5 to 4 Hz, theta: 4 to 8 Hz, alpha: 8 to 13 Hz, beta: 13 to 30 Hz, gamma: 30 to 45 Hz) across all 30 channels, yielding 150 features. Differential entropy features compute for each band-filtered signal, adding another 150 features. Alpha asymmetry features compute for six hemisphere pairs, adding 6 features that capture frontal asymmetry.
The 306-dimensional feature vector feeds a simple MLP (306128645) with batch normalization and 50% dropout. Despite having more parameters than EEGNet, this approach operates on meaningful, low-noise features rather than raw high-dimensional signals.
Illustrative Comparison.
The raw EEG segment contains 15,000 samples (30 channels 500 time points) and includes substantial nuisance variability. In contrast, the 306-dimensional feature representation applies (i) a frequency-domain transformation with averaging (Welch), (ii) band selection that emphasizes physiologically meaningful rhythms (e.g., alpha/beta), and (iii) explicit hemispheric asymmetry features. Learning an equivalent denoising and summarization pipeline end-to-end from 280 training samples per subject is unlikely to generalize.
4.3.4 M3-Vision v1: SE Block Fix
We found that the squeeze-excitation (SE) module in the original video CNN used reduction ratio 1, i.e., no bottleneck (204820482048 rather than 20481282048). Since SE relies on a bottleneck to learn channel reweighting efficiently Hu et al. (2018), we set the reduction ratio to 16.
4.3.5 M3-Vision v2: Delta Features
Following our audio success, we added temporal delta features to the video CNN, capturing frame-to-frame changes in facial expression. Delta features compute the difference between consecutive frame representations, encoding expression dynamics that static features miss.
5 Experiments and Results
All models were trained on NVIDIA RTX 6000 Ada GPUs with 48GB of VRAM. We report mean accuracy and standard deviation across all 42 subjects.
| Category | Model | EEG | Audio | Vision |
| Paper Baselines | CNN | 60.00% | 61.90% | 71.40% |
| Transformer | 53.50% | 62.70% | 74.50% | |
| M1: Our Baseline | Transformer | 52.68% 10.3 | 58.06% 10.5 | 75.30% 10.5 |
| M2: Factorized Attn. | EEG Tri-Stream | 48.00% 12.1 | — | — |
| EEGNet-Improved | 58.47% 11.2 | — | — | |
| Audio Dual-Attn. | — | 49.46% 15.8 | — | |
| Vision ViViT-style | — | — | 69.54% 12.3 | |
| M3: CNN Improved | EEG v1 (Bug Fix) | 57.66% 13.7 | — | — |
| EEG v2 (Band Power) | 67.62% 8.2 | — | — | |
| Audio (Delta MFCC) | — | 65.56% 9.3 | — | |
| Vision v1 (SE Fix) | — | — | 69.42% 11.9 | |
| Vision v2 (Delta) | — | — | 72.68% 10.8 |
| Subject | EEG | Audio | Vision |
| Top Performers | |||
| S17 | 82.50% | 85.00% | 93.33% |
| S6 | — | 78.33% | 94.17% |
| S20 | 83.33% | — | 91.67% |
| S24 | 85.00% | — | — |
| S2 | 80.83% | 85.00% | — |
| Difficult Subjects | |||
| S35 | 46.67% | 55.00% | 56.67% |
| S12 | — | 54.17% | 43.33% |
| S29 | 56.67% | 47.50% | — |
| S37 | — | 46.67% | — |
| S34 | 54.17% | 49.17% | — |
Table 1 summarizes results across all models. Three findings emerge.
M2 Factorized Attention Underperforms.
All M2 models fall below their respective baselines. The EEG Tri-Stream reaches 48% (vs. a 60% paper baseline), Audio Dual-Attention reaches 49.46% (with several subjects near chance-level performance), and the ViViT-style video model reaches 69.54% (vs. 74.5%). Although the factorized designs are well motivated, they appear difficult to fit reliably in the low-sample regime ( 280 training samples per subject).
M3 Simple Fixes Improve Performance.
Domain-appropriate feature design yields consistent gains. EEG band-power features reach 67.62% (+7.62pp over the paper baseline), Audio delta MFCCs reach 65.56% (+3.66pp), and Vision delta features reach 72.68% (+1.28pp over the paper CNN). These results suggest that, under limited supervision, established signal-processing priors (frequency analysis and temporal derivatives) can provide more reliable inductive bias than additional attention modules. Figure 2 summarizes these models.
Vision: Pretraining Domain Matters.
For Vision, M1’s domain-specific pretraining (75.30%) outperformed all M3 CNN modifications (72.68%), demonstrating that pretraining domain matters more than architectural changes. The ViT model pretrained on facial emotions provides better features than ResNet50 pretrained on ImageNet, even with our improvements.
Subject Variability.
Table 2 shows substantial individual differences. Subject 17 consistently achieves the highest accuracy across all modalities (93.33% vision, 85% audio, 82.5% EEG), suggesting relatively expressive emotional displays. Subjects 35 and 12 present persistent challenges, potentially due to subtle expressions or individual physiological differences.
6 Analysis: Why M2 Underperformed
Our M2 architectures likely underperformed for three interconnected reasons.
Degradation of Pretrained Representations.
AST and ViT backbones encode strong representations learned from large-scale pretraining. Although we include skip connections, the additional attention layers still perturb these features. The learned skip weight () indicates a strong preference for the pretrained pathway ( 88%), yet the residual attention contribution can still negatively affect downstream accuracy.
Insufficient Training Data.
The EEG Tri-Stream model has 138,401 parameters but only 280 training samples per subject. In this setting, model capacity can easily exceed what the data can constrain, leading to high variance and unstable generalization. Empirically, we observed strong overfitting (near-perfect training accuracy with substantially lower test accuracy), particularly for the M2 audio variant where multiple subjects collapsed toward chance-level performance.
Complex Inductive Biases Require Data.
Factorized attention introduces additional structure (e.g., separate spatial, temporal, and frequency reasoning). However, exploiting these inductive biases still requires sufficient supervision to estimate the extra parameters and to learn stable fusion. With limited data, the added flexibility can dominate any benefit and reduce robustness.
7 Analysis: Why M3 Succeeded
Domain Knowledge Over Learned Features.
Band power features incorporate well-established neurophysiological priors into a compact 306-dimensional representation. Alpha asymmetry directly operationalizes Davidson’s frontal asymmetry model. Delta MFCCs capture short-term speech dynamics that are commonly associated with affect. These features emphasize signal components that are difficult to learn reliably from limited data.
Implementation Fidelity Dominates in the Low-Sample Regime.
With only 280 training samples per subject, small implementation mismatches can materially change the effective learning problem and overwhelm the benefits of additional architectural components. We therefore performed targeted consistency checks that preserve the intended objective and inductive bias: (i) ensuring loss–activation compatibility by training on raw logits with CrossEntropyLoss (avoiding redundant softmax operations that distort gradients), and (ii) enforcing a true squeeze-and-excitation bottleneck via a non-trivial reduction ratio (so the module performs channel reweighting rather than a near-identity mapping). These corrections improve methodological validity and, in several cases, recover accuracy without increasing model capacity.
Minimal Complexity.
M3 models add minimal parameters over baselines. The 220-feature audio CNN adds only 40 dimensions. The band power MLP, while larger than EEGNet, operates on preprocessed features rather than raw 15,000-dimensional signals.
8 Limitations
Our study has several limitations. First, all experiments use subject-dependent evaluation; cross-subject generalization remains unexplored. Second, the EAV dataset’s conversational paradigm may not generalize to spontaneous emotions. Third, we did not explore multimodal fusion, which may further improve results by combining complementary information across modalities. Fourth, our M2 architectures may benefit from more careful hyperparameter tuning on larger validation sets. Finally, the EAV dataset contains only 42 subjects, limiting statistical power for detecting small effects.
9 Conclusion
This study systematically investigated attention mechanisms for multimodal emotion recognition on small datasets. Our findings show that for datasets with approximately 280 training samples per subject, domain-appropriate signal processing outperforms sophisticated neural architectures.
Our best results (EEG 67.62%, +7.62pp; Audio 65.56%, +3.66pp; Vision 75.30%, +0.80pp transformer and 72.68%, +1.28pp CNN) were achieved through simple modifications: frequency-domain features for EEG, delta coefficients for audio and video, and properly pretrained transformers. Meanwhile, our theoretically motivated factorized attention mechanisms (M2) consistently underperformed by 5 to 13 percentage points.
The practical lesson is clear: before adding architectural complexity, researchers should (1) verify baseline implementations for bugs, (2) consider domain-specific feature engineering, and (3) match model capacity to available training data. Future work should evaluate these factorized attention approaches on larger emotion datasets where their sophisticated inductive biases may prove beneficial.